How is pH calculated? Explain the meaning of the calculation, giving some examples of pH numbers and what they mean. How much more acidic is a solution with a pH of 3 compared to a pH of 4How is pH calculated? Explain the meaning of the calculation, giving some examples of pH numbers and what they mean. How much more acidic is a solution with a pH of 3 compared to a pH of 4.
Answer #1

The pH of any solution is measure of the concentration of hydronium ions(H3O+) present. The formula to calculate pH is -

pH = -log[H3O+]

given pH are 3 nad 4, now

3 = -log[H3O+] =[H3O+] = e-3 =4.98 * 10-2M

4 = -log[H3O+] =[H3O+] =e-4 =1.83 *10-2M

HERE SUBSTANCE WITH pH 3 HAS HIGHER CONCENTRATION OF HYDRONIUM IONS WHICH MAKES IT A STRONGER BASE.

NOTE THAT LOWER THE pH STRONGER THE ACID AND VICE VERSA.